Fermi National Laboratory


Accelerator Update for July 29, 2005 - August 1, 2005

Friday July 29
The day shift began with Operations monitoring store 4308, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using colliding beam, and with MiniBooNE, NuMI, and Meson MCenter taking beam.

At 8:38 AM, a Booster RF station (BRF19) tripped off for the first of four times in about ninety minutes. It reset okay each time.

Operations put the Recycler into shot setup at 10:37 AM to stash antiprotons. Operators completed the transfers an hour later.

The MCenter and MTest experiments began taking beam around 11:37 AM.

Pbar resumed stacking at 11:49 AM.

At 1:28 PM, BRF19 tripped off. It reset okay.

A Booster expert began a damper study at 2:03 PM. A Main Injector (MI) expert began a slipstacking study at the same time.

At 6:14 PM, BRF19 tripped off. It reset okay.

At 8:40 PM, the antiproton production and stacking rate dropped significantly. Operations investigated and called in experts.

MTest reported at 11:04 PM that they were finished with the beam for the night.

Operations reported at 11:08 PM that the Accumulator ARF1 crystal oscillator was the cause for the drop in antiproton production.

Saturday July 30
The midnight shift began with Operations monitoring store 4308 and the Recycler stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations terminated store 4308 at 2:03 AM.

Operations put the Tevatron (TeV) into shot setup at 2:33 AM.

Operations established combination store 4309 at 5:11 AM with an initial luminosity of 81.77E30. Operators reported that the first transfer of antiprotons was lost. Pbar resumed stacking.

MCenter resumed taking beam at 5:17 AM.

At 8:17 AM, BRF19 tripped off for the first of three times in about four hours.

The Recycler began stashing antiprotons at 9:45 AM. The stash ended at 10:24 AM.

Pbar resumed stacking at 10:33 AM and MCenter and MTest resumed taking beam a few minutes later.

At 1:34 PM, a Booster beam problem halted beam. Operators investigated a discovered a kicker (MKS05) problem. They called in experts.

At 4:32 PM, Operations prepared Booster for an access to repair MKS05. The experts replaced the kicker, repaired a cable termination, and were out of the tunnel by 6:30 PM.

At 7:20 PM, the Booster experts reported having trouble running beam.

At 9:20 PM, Operations turned off MI to replace a CAMAC timing card for a MI52 kicker.

A Booster expert reported at 10:29 PM that MKS05 was up and running. All downstream beam resumed by 10:47 PM.

At 11:23 PM, the TeV suffered a multi-house quench. Later it was discovered that a TeV sector B0 quadrupole power supply (B0Q6) had ramped down and caused the beam orbit to move.

Sunday July 31
The midnight shift began with Operations monitoring the stash, with the TeV recovering from a quench, with Pbar stacking, and with MiniBooNE, NuMI, MTest, and MCenter taking beam.

MTest finished using the beam at 1:27 AM.

Operations put the TeV into a dry squeeze at 5:10 AM.

Operations put the TeV into shot setup at 6:10 AM.

Operations established combination store 4312 at 9:55 AM with an initial luminosity of 76.5E30. Operators reported that there was some antiproton fallout in MI with every transfer. Pbar resumed stacking.

MCenter and MTest resumed taking beam at 10:13 AM.

A Recycler expert began a study period at 11:06 AM.

The Recycler began stashing antiprotons around 3 PM and finished around 4 PM. Pbar resumed stacking.

MCenter and MTest resumed taking beam at 4:12 PM.

The Recycler began stashing antiprotons around 8:21 PM and finished around 9:07 PM. Pbar resumed stacking.

MCenter and MTest resumed taking beam at 9:18 PM.

Monday August 1
The midnight shift began with Operations monitoring store 4312 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

At 2:59 AM, a Booster kicker MKS03 tripped off. Operators reported at 3:53 AM that the kicker was still off and that they were giving it some time to cool off. This halted Switchyard beam.

The Plans for Monday
The plan for today is to terminate store 4312 around 10:30 AM and then put a new store in.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/1/2005   email Fermilab